# Dynamical Formation of Apparent Horizons due to Boundary Effect in Vacuum Einstein Gravity

### v2

@@ 6. $L^{2}(S_{u,\underline{u}})$ Estimates for the Weyl curvature components
(155) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}(\Psi_{u},\alpha)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\lesssim\frac{\Gamma^{2}}{\lvert u\rvert}\lesssim 1.$
@@
(156) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\underline{\hat{\chi}})\nabla^{i_{4}}(\Psi_{u},\alpha)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\int_{0}^{\underline{u}}\frac{\lvert u\rvert}{a}\frac{\Gamma^{2}}{\lvert u\rvert}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim 1.$
@@
(157) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\int_{0}^{\underline{u}}\frac{a^{\frac{1}{2}}\Gamma^{2}}{\lvert u\rvert}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim 1.$
@@
(158) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\int_{0}^{\underline{u}}\frac{a^{\frac{3}{2}}\Gamma^{3}}{\lvert u\rvert^{2}}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim 1.$
@@
$\sum_{0\leq i\leq N+2}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\alpha}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim 1.$
@@
$\hskip 19.91692pt\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\alpha}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}$
$\lesssim$ $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\underline{\alpha}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
@@
In the above expression, all terms can be bounded above by $1$, in the same way as in the preceding Proposition. The result follows.
@@ 6.1. Energy Estimates for the Weyl/Riemann Curvature components
$\nabla_{3}\nabla^{i}\Psi_{1}+\left(\frac{i+1}{2}+s_{2}(\Psi_{1})\right)\mbox{tr}\underline{\chi}\nabla^{i}\Psi_{1}-\mathcal{D}\nabla^{i}\Psi_{2}=P_{i},$
$\nabla_{4}\nabla^{i}\Psi_{2}-\prescript{*}{}{\mathcal{D}}\nabla^{i}\Psi_{1}=Q_{i},$
@@
(161) $\begin{split}&\int_{H_{u}^{(0,\underline{u})}}\lVert{\nabla^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\int_{\underline{H}_{\underline{u}}^{(u_{\infty},u)}}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{\nabla^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\int_{H_{u_{\infty}}^{(0,\underline{u})}}\lVert{\nabla^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u_{\infty},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\int_{\underline{H}_{0}^{(u_{\infty},u)}}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{\nabla^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},0})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\\ &+\iint_{\mathcal{D}_{u,\underline{u}}}\frac{a}{\lvert u^{\prime}\rvert}\lVert{\nabla^{i}\Psi_{1}\cdot P_{i}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\iint_{\mathcal{D}_{u,\underline{u}}}\frac{a}{\lvert u^{\prime}\rvert}\lVert{\nabla^{i}\Psi_{2}\cdot Q_{i}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
@@
Proposition 6.5.
Let $f(x,y),\hskip 1.42262ptg(x,y)$ be positive functions defined on a rectangle $U:=\begin{Bmatrix}0\leq x\leq x_{0},\hskip 1.42262pt0\leq y\leq y_{0}\end{Bmatrix}.$ Suppose there exist constants $J,c_{1},c_{2}$ such that
$f(x,y)+g(x,y)\lesssim J+c_{1}\int_{0}^{x}f(x^{\prime},y)\hskip 1.42262pt\text{d}x^{\prime}+\int_{0}^{y}g(x,y^{\prime})\hskip 1.42262pt\text{d}y^{\prime},$
for all $(x,y)\in U$. Then there holds
$\forall(x,y)\in U:\hskip 14.22636ptf(x,y)+g(x,y)\lesssim J\mathrm{e}^{c_{1}x+c_{2}y}.$
@@
Proposition 6.6.
Under the assumptions of Theorem 3.1 and the bootstrap assumptions (102), there holds, for all $0\leq i\leq N+4$:
(162) $\begin{split}&\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}\\ &\lesssim\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}+1.\end{split}$
@@
$\nabla_{3}\nabla^{i}\alpha+\frac{i+1}{2}\mbox{tr}\underline{\chi}\nabla^{i}\alpha-\mathcal{D}\nabla^{i}\beta$
@@
$\nabla_{4}\nabla^{i}\beta-\prescript{*}{}{\mathcal{D}}\nabla^{i}\alpha=$ $\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\hat{\chi})\nabla^{i_{4}}(\beta,\alpha)$
@@
$\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{6}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\lesssim\Gamma^{3}.$
For the seventh term, there holds
@@
$\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{8}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\lesssim\Gamma^{2}+\Gamma R.$
For the ninth and tenth terms, there holds
$\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{9}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}+\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}P_{i}^{10}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\lesssim\Gamma^{3}+\Gamma^{2}R.$
@@
(167) $\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}P_{i}(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\Gamma^{3}+\Gamma^{2}R+\Gamma R+1.$
@@
(169) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a\Gamma(R+\Gamma)}{\lvert u\rvert}$
@@
(170) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a\Gamma(M+\Gamma)}{\lvert u\rvert}$
@@
(171) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{3}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}+\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{4}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\frac{a\Gamma(M+\Gamma)}{\lvert u\rvert}\end{split}$
@@
(175) $\sum_{j=1}^{7}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{j}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a\Gamma(R+M+\Gamma)}{\lvert u\rvert}+1.$

### v4

@@ 5. $L^{2}(S_{u,\underline{u}})$ Estimates for the Weyl curvature components
(189) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}=i-1}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}+1}(\Psi_{u},\alpha)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\lesssim\frac{\Gamma^{2}}{\lvert u\rvert}\lesssim a^{-\frac{1}{2}}.$
@@
(190) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\underline{\hat{\chi}})\nabla^{i_{4}}(\Psi_{u},\alpha)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\int_{0}^{\underline{u}}\frac{\lvert u\rvert}{a}\frac{\Gamma^{2}}{\lvert u\rvert}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim a^{-\frac{1}{2}}.$
@@
(191) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\int_{0}^{\underline{u}}\frac{\Gamma^{2}}{\lvert u\rvert}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim a^{-\frac{1}{2}}.$
@@
(192) $\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i-2}\nabla^{i_{1}}\psi_{g}^{i_{2}+1}\nabla^{i_{3}}(\hat{\chi},\mbox{tr}\chi)\nabla^{i_{4}}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\int_{0}^{\underline{u}}\frac{a\Gamma^{3}}{\lvert u\rvert^{2}}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim a^{-\frac{1}{2}}.$
@@
$\sum_{0\leq i\leq N+2}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\alpha}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\lesssim a^{-\frac{1}{2}}.$
@@
[…]
$+\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\underline{\alpha}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}\nabla^{i+1}\Psi_{u}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}\underline{u}^{\prime}$
@@
In the above expression, all terms can be bounded above by $a^{-\frac{1}{2}}$, in the same way as in the preceding Proposition. The result follows.
@@ 5.1. Energy Estimates for the Weyl/Riemann Curvature components
$\nabla_{3}\nabla^{i}\Psi_{1}+\left(\frac{i+1}{2}+s_{2}(\Psi_{1})\right)\mbox{tr}\underline{\chi}\nabla^{i}\Psi_{1}-D\nabla^{i}\Psi_{2}=P_{i},$
$\nabla_{4}\nabla^{i}\Psi_{2}-\prescript{*}{}{D}\nabla^{i}\Psi_{1}=Q_{i},$
@@
(195) $\begin{split}&\int_{H_{u}^{(0,\underline{u})}}\lVert{\nabla^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\int_{\underline{H}_{\underline{u}}^{(u_{\infty},u)}}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{\nabla^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\int_{H_{u_{\infty}}^{(0,\underline{u})}}\lVert{\nabla^{i}\Psi_{1}}\rVert_{L^{2}_{(sc)}(S_{u_{\infty},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\int_{\underline{H}_{0}^{(u_{\infty},u)}}\frac{a}{\lvert u^{\prime}\rvert^{2}}\lVert{\nabla^{i}\Psi_{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},0})}^{2}\hskip 1.42262pt\text{d}u^{\prime}\\ &+\iint_{D_{u,\underline{u}}}\frac{a}{\lvert u^{\prime}\rvert}\lVert{\nabla^{i}\Psi_{1}\cdot P_{i}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}+\iint_{D_{u,\underline{u}}}\frac{a}{\lvert u^{\prime}\rvert}\lVert{\nabla^{i}\Psi_{2}\cdot Q_{i}}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}.\end{split}$
@@
Proposition 5.5 (Two–parameter Grönwall inequality on a rectangle).
Let $x_{0},y_{0}>0$ and set
$U:=\{(x,y)\in\mathbb{R}^{2}:\;0\leq x\leq x_{0},\;0\leq y\leq y_{0}\}.$
Let $f,g:U\to[0,\infty)$ be measurable functions such that for each fixed $y$ the map $x\mapsto f(x,y)$ is integrable on $0,x_{0}]$ and for each fixed $x$ the map $y\mapsto g(x,y)$ is integrable on $0,y_{0}]$. Assume there exist constants $J,c_{1},c_{2}\geq 0$ and $C_{0}\geq 1$ such that for all $(x,y)\in U$ one has
(196) $f(x,y)+g(x,y)\leq C_{0}\!\left(J+c_{1}\int_{0}^{x}f(x^{\prime},y)\,dx^{\prime}+c_{2}\int_{0}^{y}g(x,y^{\prime})\,dy^{\prime}\right).$
Then for every $(x,y)\in U$ there holds the exponential bound
(197) $f(x,y)+g(x,y)\leq C\,J\,\exp\!\big(C(c_{1}x+c_{2}y)\big),$
where $C\geq 1$ depends only on $C_{0}$ (in particular, it is independent of $x_{0},y_{0},J,c_{1},c_{2}$ and of the functions $f,g$).
@@
Proposition 5.6 (Top–order curvature flux estimate for $(\alpha,\beta)$).
Let $(\mathcal{M},g)$ be a smooth vacuum spacetime endowed with a regular double–null foliation $(u,\underline{u})$ and associated null frame $(e_{3},e_{4},e_{A})$. Assume the hypotheses of Theorem 2.2 and the bootstrap bounds (118) hold on the spacetime region under consideration. Let $N\geq 0$ denote the maximal derivative level in the energy hierarchy.
Then for every integer $0\leq i\leq N+4$, the following renormalized curvature flux estimate holds for the extreme null Weyl component $\alpha$ and the adjacent component $\beta$:
(198) $a^{\frac{1}{2}}\,\lVert{\nabla^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u}^{(0,\underline{u})})}\;+\;a^{\frac{1}{2}}\,\lVert{\nabla^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}\;\lesssim\;a^{\frac{1}{2}}\,\lVert{\nabla^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}\;+\;a^{\frac{1}{2}}\,\lVert{\nabla^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}\;+\;1.$
The implicit constant depends only on $N$, and is uniform in $a$, $u$, and $\underline{u}$.
In particular, the scale–renormalized outgoing flux of $\alpha$ along $H_{u}$ together with the incoming flux of $\beta$ along $\underline{H}_{\underline{u}}$ are controlled solely by the corresponding initial null fluxes on $H_{u_{\infty}}$ and $\underline{H}_{0}$, up to a universal error bound generated by lower–order curvature components and Ricci coefficients.
@@
$\nabla_{3}\nabla^{i}\alpha+\frac{i+1}{2}\mbox{tr}\underline{\chi}\nabla^{i}\alpha-D\nabla^{i}\beta$
@@
$\nabla_{4}\nabla^{i}\beta-\prescript{*}{}{D}\nabla^{i}\alpha=$ $\sum_{i_{1}+i_{2}+i_{3}+i_{4}=i}\nabla^{i_{1}}\psi_{g}^{i_{2}}\nabla^{i_{3}}(\psi_{g},\hat{\chi})\nabla^{i_{4}}(\beta,\alpha)$
@@
(203) $\int_{0}^{\underline{u}}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert}\lVert{a^{\frac{i}{2}}P_{i}(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{1}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}\hskip 1.42262pt\text{d}u^{\prime}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\lesssim\frac{a^{\frac{1}{2}}}{|u|}\bigg(\Gamma^{3}+\Gamma^{2}R+\Gamma R+1\bigg).$
@@
(205) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{1}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}\Gamma(R+\Gamma)}{\lvert u\rvert}$
@@
(206) $\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{2}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}\Gamma(M+\Gamma)}{\lvert u\rvert}$
@@
(207) $\begin{split}&\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{3}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}+\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{4}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\\ \lesssim&\frac{a^{\frac{1}{2}}\Gamma(M+\Gamma)}{\lvert u\rvert}\end{split}$
@@
(211) $\sum_{j=1}^{7}\int_{u_{\infty}}^{u}\frac{a}{\lvert u^{\prime}\rvert^{2}}\left(\int_{0}^{\underline{u}}\lVert{a^{\frac{i}{2}}Q_{i}^{j}}\rVert_{L^{2}_{(sc)}(S_{u^{\prime},\underline{u}^{\prime}})}^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\right)^{\frac{1}{2}}\hskip 1.42262pt\text{d}u^{\prime}\lesssim\frac{a^{\frac{1}{2}}}{|u|}\bigg((\Gamma(R+M+\Gamma)+1\bigg).$